# Representation Ensembling for Synergistic Lifelong Learning with Quasilinear Complexity

### v18

@@ 3.1.2 Synergistic Forests
Synergistic Forests (SynF) ensemble decision trees or forests. For each task, the encoder $u_{t}$ of SynF is the representation learned by a decision forest [41, 38]. The leaf nodes of each decision forest partition the input space $\mathcal{X}$ into polytopes [33]. The channel then learns the class-conditional posteriors by populating the polytopes with out-of-task samples, as in “honest trees” [33, 34, 35]. Each channel outputs the posteriors averaged across the collection of forests learned over different tasks. The decoder $w_{t}$ outputs the argmax to produce a single prediction.
@@
Note that the amount of additional representation capacity added per task by SynF is a function of the amount and complexity of the data for a new task. Contrast this with SynN  and other deep net based modular or representation ensembling approaches, which a priori choose how much additional representation to add, prior to seeing all the new task data. So, SynF has capacity, space complexity, and time complexity scale with the complexity and sample size of each task. In contrast, ProgNN, SynN (and others like it) have a fixed capacity for each task, even if the tasks have very different sample sizes and complexities.
@@ 4 A computational taxonomy of lifelong learners
Figure 3: Storage space as a function on number of tasks in CIFAR 10X10. Memory consumed by SynN is dominated by the encoder size. The size of DF-CNN remains constant throughout.
@@
    The number of epochs is fixed for each task, independent of sample size.
@@
semi-parametric $T$ $T+n$ $n$ $n$ $n$ SynN, Model Zoo, DER, LMC
non-parametric $n$ $n$ $n$ $n$ $n$ SynF, IBP-WF
@@
Assumption $3$ enables us to write time complexity as a function of the sample size. Table 1 summarizes the capacity, space and time complexity of several reference algorithms, as well as our SynN and SynF. For space and time complexity, the table shows results as a function of $n$ and $T$, as well as the common scenario where sample size per task is fixed and therefore proportional to the number of tasks, $n\propto T$. For detailed calculation of time complexity see Appendix E.
@@
Parametric lifelong learning methods have a representational capacity which is invariant to sample size and task number. Although the space complexity of some of these algorithms grow (because the number of the constraints grows stored by the algorithms grows, or they continue to store more data), their capacity is fixed. Thus, given a sufficiently large number of tasks, in general, eventually all parametric methods will catastrophically forget. EWC [10], Online EWC [13], SI [11], and LwF [12] are all examples of parametric lifelong learning algorithms.
@@
Semi-parametric algorithms’ representational capacity grows slower than sample size. For example, if $T$ is increasing slower than $n$ (e.g., $T\propto\log n$), then algorithms whose capacity is proportional to $T$ are semi-parametric. ProgNN [16] is semi-parametric, nonetheless, its space complexity $\tilde{\mathcal{O}}(T^{2})$ due to the lateral connections. Moreover, the time complexity for ProgNN also scales quadratically with $n$ when $n\propto T$. Thus, an algorithm that literally stores all the data it has ever seen, and retrains a fixed size network on all those data with the arrival of each new task, would have smaller space complexity and the same time complexity as ProgNN. For comparison, we implement such an algorithm and refer to it as Total Replay. DF-CNN [17] improves upon ProgNN by introducing a “knowledge base” with lateral connections to each new column, thereby avoiding all pairwise connections. Because these semi-parametric methods have a fixed representational capacity per task, they will either lack the representation capacity to perform well given sufficiently complex tasks, and/or will waste resources for very simple tasks. SynN and SynF eliminate the lateral connections between columns of the network, thereby reducing space complexity down to $\tilde{\mathcal{O}}(T)$. Moreover, as shown in Figure 3, memory consumed by new channels is negligible compared to that of memory required for storing the encoders. Note that the time required for updating channels is negligible in comparison with the time required for training a new encoder.
Non-parametric algorithms’ representational capacity grows linearly with sample size. SynF  is a non-parametric lifelong learning algorithm with its capacity, space and time complexity all $\tilde{\mathcal{O}}(n)$, meaning that its representational capacity naturally increases with the complexity of each task. Apart from SynF, Indian Buffet Process for Weight Factors (IBP-WF) [45] proposed the only other non-parametric lifelong learning algorithm to our knowledge.
5 Providing intuition of synergistic learning through simulations
@@
Figure 4: Synergistic Forest and Synergistic Network demonstrate forward and backward transfer. (A) 750 samples from: (Ai) Gaussian XOR, (Aii) XNOR, which has the same optimal discriminant boundary as XOR, and (Aiii) R-XOR, which has a discriminant boundary that is uninformative, and therefore adversarial, to XOR. (Bi) Generalization error for XOR, and (Bii) XNOR. SynF (SynN) outperforms RF (DN) on XOR when XNOR data is available, and on XNOR when XOR data are available. (Biii) Forward and backward transfer of SynF are positive for all sample sizes, and are negative for all sample sizes for RF. Forward and backward transfer for SynN is higher than DN for all sample sizes. (Ci) In an adversarial task setting, SynF and SynN gracefully forgets XOR, whereas RFand DN  catastrophically forget and interfere. (Cii) $\mathsf{Backward~{}Transfer}$ is positive with respect to XOR when the optimal decision boundary of $\theta$-XOR is similar to that of XOR (e.g. angles far from $45^{\circ}$), and negative otherwise.. (Ciii) $\mathsf{Backward~{}Transfer}$ is a nonlinear function of the source training sample size (XOR sample size is fixed at $500$).
@@ 5.1 Forward and backward transfer in a simple environment
Consider a very simple two-task environment: Gaussian XOR and Gaussian Exclusive NOR (XNOR) (Figure 4A, see Appendix F for details). The two tasks share the exact same discriminant boundaries: the coordinate axes. Thus, transferring from one task to the other merely requires learning a bit flip of the class labels. We sample a total 750 samples from XOR, followed by another 750 samples from XNOR.
SynF and random forests (RF) achieve the same generalization error on XOR when training with XOR data (Figure 4Bi). But because RF does not account for a change in task, when XNOR data appear, RF performance on XOR deteriorates (it catastrophically forgets). In contrast, SynF continues to improve on XOR given XNOR data, demonstrating backward transfer. Now consider the generalization error on XNOR (Figure 4Bii). Both SynF and RF are at chance levels for XNOR when only XOR data are available. When XNOR data are available, RF must unlearn everything it learned from the XOR data, and thus its performance on XNOR starts out nearly maximally inaccurate, and quickly improves. On the other hand, because SynF can leverage the encoder learned using the XOR data, upon getting any XNOR data, it immediately performs quite well, and then continues to improve with further XNOR data, demonstrating forward transfer (Figure 4Biii). SynF demonstrates positive forward and backward transfer for all sample sizes, whereas RF fails to demonstrate forward or backward transfer, and eventually catastrophically forgets the previous tasks. Results for SynN and DN are qualitatively similar to those of SynF and RF  respectively.
@@ 5.2 Forward and backward transfer for adversarial tasks
In the context of synergistic learning, we informally define a task $t$ to be adversarial with respect to task $t^{\prime}$ if the true joint distribution of task $t$, without any domain adaptation, impedes performance on task $t^{\prime}$. In other words, training data from task $t$ can only add noise, rather than signal, for task $t^{\prime}$. An adversarial task for Gaussian XOR is Gaussian XOR rotated by $45^{\circ}$ (R-XOR) (Figure 4Aiii). Training on R-XOR therefore impedes the performance of SynF and SynN  on XOR, and thus backward transfer becomes negative, demonstrating graceful forgetting [46] (Figure 4Ci). Because R-XOR is more difficult than XOR for SynF (because the discriminant boundaries are oblique [47]), and because the discriminant boundaries are learned imperfectly with finite data, data from XOR can actually improve performance on R-XOR, and thus forward transfer is positive. In contrast, both forward and backward transfer are negative for RF and DN.
To further investigate this relationship, we design a suite of R-XOR examples, generalizing R-XOR from only $45^{\circ}$ to any rotation angle between $0^{\circ}$ and $90^{\circ}$, sampling 100 points from XOR, and another 100 from each R-XOR (Figure 4Cii). As the angle increases from $0^{\circ}$ to $45^{\circ}$, $\mathsf{Backward~{}Transfer}$ flips from positive ($\approx 0.18$) to negative ($\approx-0.11$) for SynF. A similar trend is also visible for SynN. The $45^{\circ}$-XOR is the maximally adversarial R-XOR. Thus, as the angle further increases, $\mathsf{Backward~{}Transfer}$ increases back up to $\approx 0.18$ at $90^{\circ}$, which has an identical discriminant boundary to XOR. Moreover, when $\theta$ is fixed at $25^{\circ}$, $\mathsf{Backward~{}Transfer}$ increases at different rates for different sample sizes of the source task (Figure 4Ciii).
Together, these experiments indicate that the amount of transfer can be a complicated function of (i) the difficulty of learning good representations for each task, (ii) the relationship between the two tasks, and (iii) the sample size of each. Appendix F further investigates this phenomenon in a multi-spiral environment.
@@ 6 Benchmark data experiments
For benchmark data, we build SynN encoders using the network architecture described in van de Ven et al. [48]. We use the same network architecture for all the benchmarking models. For the following experiments, we consider two modalities of real data: vision and language.
@@ 6.1 Reference algorithms
We compared our approaches to $15$ reference lifelong learning methods. Among them five are resource growing as well as modular approach: ProgNN [16], DF-CNN [17], LMC [49], Model Zoo[50], CoSCL [51]. Other reference algorithms are resource constrained: Elastic Weight Consolidation (EWC) [10], Online-EWC (O-EWC) [13], Synaptic Intelligence (SI) [11], Learning without Forgetting (LwF) [12], and “None”.
We also compare two variants of exact replay (Total Replay and Partial Replay) using the code provided by van de Ven et al. [48]. Both Total and Partial Replay store all the data they have ever seen, but Total Replay replays all of it upon acquiring a new task, whereas Partial Replay replays $M$ samples, randomly sampled from the entire corpus, whenever we acquire a new task with $M$ samples. Additionally, we have compared our approach with more constrained ways of replaying old task data, including Averaged Gradient Episodic Memory (A-GEM) [52], Experience Replay (ER) [53] and Task-based Accumulated Gradients (TAG) [54].
For the baseline “None”, the network was incrementally trained on all tasks in the standard way while always only using the data from the current task. The implementations for all of the algorithms are adapted from open source codes [17, 55]; for implementation details, see Appendix D.
6.2 Exploring and explaining transfer capabilities via the CIFAR 10x10 dataset
The CIFAR 100 challenge [56], consists of 50,000 training and 10,000 test samples, each a 32x32 RGB image of a common object, from one of 100 possible classes, such as apples and bicycles. CIFAR 10x10 divides these data into 10 tasks, each with 10 classes [17] (see Appendix G for details).
@@ 6.2.1 Resource growing experiments
SynF, SynN, and Model Zoo demonstrate positive forward and backward transfer for every task in CIFAR 10x10, in contrast, other algorithms do not exhibit any positive backward transfer (Figure 1 first column). Moreover, they retained their accuracy while improving transfer (Figure 1, bottom row). ProgNN had a similar degree of forward transfer, but zero backward transfer, and requires quadratic space and time in sample size, unlike SynF, SynN, and Model Zoo which all require quasilinear space and time.
Figure 5: Ablation experiments on SynN and SynF using CIFAR 10X10. (A, B) Comparison of Resource Constrained SynF with algorithms having a fixed amount of resources. SynF is the only approach that demonstrate forward (A) and backward transfer (B). (C) Building and recycling ensembles are two boundaries of a continuum, with hybrid models in the middle. SynF achieves lower (better) generalization error than other approaches until 5,000 training samples on the new task are available, but eventually a hybrid approach wins. (D, E, F) Controlled replay experiment on CIFAR 10x10. Fraction of total samples per task replayed is mentioned in parenthesis in the middle and the right plot. Left: forward transfer is invariant to replay. Middle: Backward transfer increases as more samples are replayed from the old tasks. Right: Overall transfer increases with amount of replay.
@@ 6.2.2 Ablation Experiments
Our proposed algorithms can improve performance on all the tasks (past and future) by both growing additional resources and replaying data from the past tasks. Below we do three ablation experiments using CIFAR 10X10 to measure the relative contribution of resource growth and replay on the performance of our proposed algorithms.
Resource constrained experiments
In this experiment, we devised a “resource constrained” variant of SynF experiments to observe the effect of ablating resource growth on SynF. In this constrained variant, we compare the lifelong learning algorithm to its single task variant, but ensure that they both have the same amount of resources. For example, on Task 2, we would compare SynF with 20 trees (10 trained on 500 samples from Task 1, and another 10 trained on 500 samples from Task 2) to RF with 20 trees (all trained on 500 samples Task 2). Although ablating the resource growth results in lower forward transfer compared to that of its resource growing variant (Figure 1 first column and Figure 5 top left), forward transfer remains positive after enough tasks, and backward transfer is actually invariant to this change (Figure 5, top left and center). In contrast, all of the reference algorithms that have fixed resources exhibit negative forward and backward transfer. Note that in this experiment, building the single task learners actually requires substantially more resources, specifically, $10+20+\cdots+100=550$ trees, as compared with only $100$ trees in the prior experiments. In general, to ensure single task learners use the same amount of resources per task as our synergistic learners requires $O(n^{2})$ resources, where as SynF only requires $O(n)$.
Resource Recycling Experiments
The binary distinction we made above, algorithms either build resources or reallocate them, is a false dichotomy, and biologically unnatural. In biological learning, systems develop from building to fixed resources, as they grow from juveniles to adults. To explore this continuum of amount of resources to grow, we trained SynF on the first nine CIFAR 10x10 tasks using 50 trees per task, with 500 samples per task. For the tenth task, we could (i) select the 50 trees (out of the 450 existing trees) that perform best on task 10 (recycling), (ii) train 50 new trees, as SynF would normally do (building), (iii) build 25 and recruit 25 trees (hybrid), or (iv) ignore all prior trees (RF). SynF outperforms other approaches except when 5,000 training samples are available, but the recycling approach is nearly as good as SynF (Figure 5, top right). This result motivates future work to investigate optimal strategies for determining how to optimally leverage existing resources without growing new ones given a new task.
@@
In this experiment, we train $4$ different versions of SynN and SynF sequentially on the $10$ tasks from CIFAR 10X10. The only difference between different versions of the algorithms is the amount of old task data replayed. In $4$ different versions of each algorithm, we replay $40\%$, $60\%$, $80\%$ and $100\%$ of the old task data respectively. As apparent from Figure 5 bottom, replaying old task data has no effect on forward transfer, but replaying more data improves backward transfer as the number of tasks increases.

### v19

@@ 3.1.3 Additional realization of our approach using random forest as encoder
Simple Lifelong Learning Forest (SiLLy-F) ensembles decision trees or forests. For each task, the encoder $u_{t}$ of SiLLy-F is the representation learned by a decision forest [40, 38]. The channel then learns the class-conditional posteriors by populating the forest leaves with out-of-task samples, as in “honest trees” [33, 34, 35]. Each channel outputs the posteriors averaged across the collection of forests learned over different tasks. The decoder $w_{t}$ outputs the argmax to produce a single prediction.
@@
Note that the amount of additional representation capacity added per task by SiLLy-F is a function of the amount and complexity of the data for a new task. Contrast this with SiLLy-N  and other deep net based modular or representation ensembling approaches, which a priori choose how much additional representation to add, prior to seeing all the new task data. So, SiLLy-F has capacity, space complexity, and time complexity scale with the complexity and sample size of each task. In contrast, ProgNN, SiLLy-N (and others like it) have a fixed capacity for each task, even if the tasks have very different sample sizes and complexities.
@@ 4 A computational taxonomy of lifelong learners
Figure 3: Storage space as a function on number of tasks in CIFAR 10X10. Memory consumed by SiLLy-N is dominated by the encoder size. The size of DF-CNN remains constant throughout.
@@
    The number of epochs is fixed for each task.
@@
semi-parametric $T$ $T+n$ $n$ $n$ $n$ SiLLy-N, Model Zoo, DER, LMC
non-parametric $n$ $n$ $n$ $n$ $n$ SiLLy-F, IBP-WF
@@
Assumption $3$ enables us to write time complexity as a function of the sample size. Table 1 summarizes the capacity, space and time complexity of several reference algorithms, as well as our SiLLy-N and SiLLy-F. For space and time complexity, the table shows results as a function of $n$ and $T$, as well as the common scenario where sample size is fixed per task, and therefore proportional to the number of tasks, $n\propto T$. For detailed calculation of time complexity see Appendix A.3.
@@
Parametric lifelong learning methods have a representational capacity which is invariant to sample size and task number. Although the space complexity of some of these algorithms grow (because the number of the constraints stored by the algorithms grows, or they continue to store more data), their capacity is fixed. Thus, given a sufficiently large number of tasks with increasing complexity, in general, eventually all parametric methods will catastrophically forget. EWC [10], Online EWC [13], SI [11], and LwF [12] are all examples of parametric lifelong learning algorithms. Our fixed resource algorithms are also parametric. For comparison, we implement another baseline algorithm and refer to it as Total Replay, which is also parametric. Total Replay replays both old and current task data while learning a new task.
@@
Semi-parametric algorithms’ representational capacity grows slower than sample size. For example, if $T$ is increasing slower than $n$ (e.g., $T\propto\log n$), then algorithms whose capacity is proportional to $T$ are semi-parametric. ProgNN [16] is semi-parametric, nonetheless, its space complexity is $\tilde{\mathcal{O}}(T^{2})$ due to the lateral connections. Moreover, the time complexity for ProgNN also scales quadratically with $n$ when $n\propto T$. Thus, an algorithm that literally stores all the data it has ever seen, and retrains a fixed size network on all those data with the arrival of each new task, would have smaller space complexity and the same time complexity as ProgNN. DF-CNN [17] improves upon ProgNN by introducing a “knowledge base” with lateral connections to each new column, thereby avoiding all pairwise connections. Because these semi-parametric methods have a fixed representational capacity per task, they will either lack the representation capacity to perform well given sufficiently complex tasks, and/or will waste resources for very simple tasks.
SiLLy-N and SiLLy-F eliminate the lateral connections between columns of the network, thereby reducing space complexity down to $\tilde{\mathcal{O}}(T)$. Moreover, as shown in Figure 3, memory consumed by new channels is negligible compared to that of memory required for storing the encoders. Note that the time required for updating channels is negligible in comparison with the time required for training a new encoder.
Indian Buffet Process for Weight Factors (IBP-WF) is another non-parametric lifelong learning algorithm.
5 Providing intuition of simple lifelong learning through simulations
@@
Figure 4: SiLLy-N demonstrates forward and backward transfer. (A) 750 samples from: (Ai) Gaussian XOR, (Aii) XNOR, which has the same optimal discriminant boundary as XOR, and (Aiii) R-XOR, which has a discriminant boundary that is uninformative, and therefore adversarial, to XOR. (Bi) Generalization error for XOR, and (Bii) XNOR. SiLLy-N outperforms DN on XOR when XNOR data is available, and on XNOR when XOR data are available. (Biii) Forward and backward transfer of SiLLy-N are positive for all sample sizes. (Ci) In an adversarial task setting, SiLLy-N gracefully forgets XOR, whereas DN  catastrophically forget and interfere. (Cii) $\mathsf{Backward~{}Transfer}$ is maximum positive with respect to XOR when the optimal decision boundary of $\theta$-XOR is similar to that of XOR (e.g. angles far from $45^{\circ}$), and negative otherwise. The dashed line shows the regression line fitted through the original points. (Ciii) $\mathsf{Backward~{}Transfer}$ is a nonlinear function of the source training sample size (XOR sample size is fixed at $500$).
@@ 5.1 Forward and backward transfer in a simple environment
Consider a very simple two-task environment: Gaussian XOR and Gaussian Exclusive NOR (XNOR) (Figure 4A, see Appendix A.4 for details). The two tasks share the exact same discriminant boundaries: the coordinate axes. Thus, transferring from one task to the other merely requires learning a bit flip of the class labels. We sample a total 750 samples from XOR, followed by another 750 samples from XNOR.
SiLLy-N and deep network (DN) achieve the same generalization error on XOR when training with XOR data (Figure 4Bi). But because DN does not account for a change in task, when XNOR data appear, DN performance on XOR deteriorates (it catastrophically forgets). In contrast, SiLLy-N continues to improve on XOR given XNOR data, demonstrating backward transfer. Now consider the generalization error on XNOR (Figure 4Bii). Both SiLLy-N and DN are at chance levels for XNOR when only XOR data are available. When XNOR data are available, DN must unlearn everything it learned from the XOR data, and thus its performance on XNOR starts out nearly maximally inaccurate, and quickly improves. On the other hand, because SiLLy-N can leverage the encoder learned using the XOR data, upon getting any XNOR data, it immediately performs quite well, and then continues to improve with further XNOR data, demonstrating forward transfer (Figure 4Biii). SiLLy-N demonstrates positive forward and backward transfer for all sample sizes, whereas DN fails to demonstrate neither forward nor backward transfer, and eventually catastrophically forgets the previous tasks.
@@ 5.2 Forward and backward transfer for adversarial tasks
In the context of lifelong learning, we informally define a task $t$ to be adversarial with respect to task $t^{\prime}$ if the true joint distribution of task $t$, without any domain adaptation, impedes performance on task $t^{\prime}$. In other words, training data from task $t$ can only add noise, rather than signal, for task $t^{\prime}$. An adversarial task for Gaussian XOR is Gaussian XOR rotated by $45^{\circ}$ (R-XOR) (Figure 4Aiii). Training on R-XOR therefore impedes the performance of SiLLy-N  on XOR, and thus backward transfer becomes negative, demonstrating graceful forgetting [44] (Figure 4Ci).
To further investigate this relationship, we design a suite of R-XOR examples, generalizing R-XOR from only $45^{\circ}$ to any rotation angle between $0^{\circ}$ and $90^{\circ}$, sampling 100 points from XOR, and another 100 from each R-XOR (Figure 4Cii). Note that we could not run the experiment for a lot of Monte Carlo repetition to have a smooth curve and hence we have shown a regressed curve fitted to the low repetition noisy curve. As the angle increases from $0^{\circ}$ to $45^{\circ}$, $\mathsf{Backward~{}Transfer}$ gradually decreases for SiLLy-N. The $45^{\circ}$-XOR is the maximally adversarial R-XOR. Thus, as the angle further increases, $\mathsf{Backward~{}Transfer}$ increases back up to $\approx 0.18$ at $90^{\circ}$, which has an identical discriminant boundary to XOR. Moreover, when $\theta$ is fixed at $25^{\circ}$, $\mathsf{Backward~{}Transfer}$ increases at different rates for different sample sizes of the source task (Figure 4Ciii).
Together, these experiments indicate that the amount of transfer can be a complicated function of (i) the difficulty of learning good representations for each task, (ii) the relationship between the two tasks, and (iii) the sample size of each.
@@ 6 Benchmark data experiments
For benchmark data, we build SiLLy-N encoders using the network architecture described in [45]. We use the same network architecture for all the benchmarking models. For the following experiments, we consider two modalities of real data: vision and language.
@@ 6.1 Reference algorithms
We compared our approaches to $15$ reference lifelong learning methods. Among them five are resource growing as well as modular approach: ProgNN [16], DF-CNN [17], LMC [46], Model Zoo[47], CoSCL [48]. Note that "Model Zoo" was published after our work was archived on arXiv, and the authors have built on our work (personal communications). Other reference algorithms are resource constrained: Elastic Weight Consolidation (EWC) [10], Online-EWC (O-EWC) [13], Synaptic Intelligence (SI) [11], Learning without Forgetting (LwF) [12], and “None”.
We also compare two variants of exact replay (Total Replay and Partial Replay) using the code provided in [45]. Both Total and Partial Replay store all the data they have ever seen, but Total Replay replays all of it upon acquiring a new task, whereas Partial Replay replays $N$ samples, randomly sampled from the entire corpus, whenever we acquire a new task with $N$ samples. Additionally, we have compared our approach with more constrained ways of replaying old task data, including Averaged Gradient Episodic Memory (A-GEM) [49], Experience Replay (ER) [50] and Task-based Accumulated Gradients (TAG) [51].
For the baseline “None”, the network was incrementally trained on all tasks in the standard way while always only using the data from the current task. The implementations for all of the algorithms are adapted from open source codes [17, 52]; for implementation details, see Appendix A.2.
6.2 Exploring and explaining transfer capabilities via CIFAR 10x10 dataset
The CIFAR 100 challenge [53], consists of 50,000 training and 10,000 test samples, each a 32x32 RGB image of a common object, from one of 100 possible classes, such as apples and bicycles. CIFAR 10x10 divides these data into 10 tasks, each with 10 classes [17] (see Appendix A.5 for details).
@@ 6.2.1 Resource growing experiments
SiLLy-N and Model Zoo demonstrate positive forward and backward transfer for every task in CIFAR 10x10, in contrast, other algorithms do not exhibit any positive backward transfer (Figure 1 first column). Moreover, they retained their accuracy while improving transfer (Figure 1, bottom row). ProgNN had a similar degree of forward transfer, but zero backward transfer, and requires quadratic space and time in sample size, unlike SiLLy-N which requires quasilinear space and time.
Figure 5: Ablation experiments on SiLLy-N using CIFAR 10X10. Top row: SiLLy-N uniquely shows both positive forward and backward transfer while operating with the same number of parameters as other baseline constant parameter approaches. Bottom row: Replaying old task data impacts backward transfer while keeping forward transfer unchanged for SiLLy-N.
@@ 6.2.2 Ablation experiments
Our proposed algorithms can improve performance on all the tasks (past and future) by both growing additional resources and replaying data from the past tasks. Below we do two ablation experiments using CIFAR 10X10 to measure the relative contribution of resource growth and replay on the performance of our proposed algorithms.
Constrained resource experiment
In this experiment, we ablate the capability of SiLLy-N to grow additional resources after learning $4$ encoders. We also reduce the number of channels and nodes at each encoder layer by four times to keep the total number of parameters similar to the other constant-resource-algorithms. As shown in the top row of Figure 5, SiLLy-N-4 still shows positive forward and backward transfer with constant resources. However, the accuracy for SiLLy-N-4 gets reduced compared to that of resource growing SiLLy-N in Figure 1. Note that all the baseline algorithms have negative backward transfer. This experiment indicates that constant resource mode operation for SiLLy-N may be advantageous when we have a lot of tasks to learn and have a decent amount of storage budget available. We will elaborate the above point later with a large scale dataset (food1k).
@@
In this experiment, we train four different versions of SiLLy-N sequentially on the $10$ tasks from CIFAR 10X10. The only difference between different versions of the algorithms is the amount of old task data replayed. In four different versions of each algorithm, we replay $40\%$, $60\%$, $80\%$ and $100\%$ of the old task data respectively. As apparent from Figure 5 bottom, replaying old task data has no effect on forward transfer, but replaying more data improves backward transfer as the number of tasks increases.
6.2.3 Experiment using pretrained encoders
Figure 6: Pretrained encoder on CIFAR 10X10. Using pretrained encoders results in better transfer and accuracy for SiLLy-N.
In this experiment, we explore the effect of using pretrained encoders on the performance of SiLLy-N. For this experiment only, we use ResNet 50, as weights for the model pretrained on imagenet dataset is publicly available with the python package: Keras [54]. We initialize the weights of the encoders with the pretrained weights and do not freeze any layer during training. As shown in Figure 6, pretraining the encoders results in better accuracy and forward transfer, but less backward transfer for SiLLy-N. However, as it is unclear how to use pretrained encoders for other baseline approaches, we do not use pretrained encoders for SiLLy-N in other experiments for fair comparison.